# From Unimodal to Multimodal: Scaling up Projectors to Align Modalities

### v1

@@ 4.2 Impact of Projector Architectures
We ablate our projector combinations for the DINOv2 and All-Roberta-Large-v1 encoders by training the projectors to convergence on the LAION-Class-Collected dataset and evaluating the performance on ImageNet 0-shot domain transfer. An MLP applied solely to the local vision tokens achieved 68.81% accuracy, while a Token projection Mukhoti et al. (2023) performed slightly better. Therefore, we used the Token projector for all tokens, both visual and textual. Adding projectors to the text side, targeting both text tokens and a global projector on the averaged local tokens (rows 3, 4, and 5), resulted in performance improvements. These projectors help transform the unimodal text encoder’s language-only representations to be more similar to the visual representations. Introducing projectors to the CLS token (row 6) of the visual encoder led to a significant performance increase from 72.15% to 75.13%. Using both CLS and patch projectors in tandem yielded the best performance at 76.12%. This improvement is attributed to DINOv2’s dual training objectives: the image-level DINO Caron et al. (2021) objective on the CLS token and the patch-level iBOT Zhou et al. (2021) objective on the patch tokens learning effective global and local features.
@@ 4.3 Impact of Class-Collected Data / Retrieval Data
In this section, we ablate the different components of our alignment data. Specifically, we compare the high concept coverage LAION-CLASS-Collected dataset with the higher image-caption quality retrieval datasets: CC3M, CC12M, and SBU. Our experiments show that aligning DINOv2 and All-Roberta-Large-v1 on the high concept coverage dataset results in a high ImageNet zero-shot domain transfer accuracy of 76.1 , though the retrieval accuracies are lower, at 52.7/42.2. In contrast, training with the higher image-caption quality retrieval datasets results in high image and text retrieval scores on the Flickr30k val set (85.3 and 72.4, respectively). However, the limited concept coverage of these datasets leads to a much lower ImageNet accuracy of 54.1. Combining both types of datasets yields both high ImageNet accuracy and high image/text retrieval accuracies. To ensure that the extra data is adequately utilized, we train for an additional 30 epochs. This approach results in our best-performing model, achieving an ImageNet accuracy of 76.30 and Flickr retrieval scores of 87.54/74.17 (last row).
@@ 5 Results
We evaluate the alignment between vision and text encoders across benchmarks commonly used for CLIP-like models, including zero-shot image classification, image retrieval, localization, multilingual classification/retrieval, and dense caption image-text retrieval. We demonstrate that aligning unimodal vision-language encoders can match or exceed the performance of large CLIP models, despite using smaller datasets and less compute. Additionally, our alignment framework is flexible, enabling the use of specialized encoders for specific tasks, such as aligning multilingual text encoders for multilingual or low-resource image classification/retrieval, or long-context text encoders for dense image/caption retrieval. Furthermore, aligning DINOv2 with a text encoder improves image localization beyond CLIP’s vision encoder due to DINOv2’s superior localization features.
@@
Tables 3 and 4 report our model’s performance on zero-shot domain transfer to image classification datasets and zero-shot image-text retrieval on the Flickr30k and COCO datasets, respectively. Similar to Maniparambil et al. (2023) we use classwise Visually Descriptive Text (VDT) prompts to enable the unimodal-text encoder in our DINOv2-ARL projector model to better identify the zero-shot classes of the downstream datasets. Detailed descriptions of the evaluation datasets can be found in the A.10, highlighting dataset domains, sizes, and prompt descriptions. We see that despite being trained on a 20M dataset our DINOv2-ARL projector model achieves an ImageNet accuracy of 76.3 which is 1 and 3.6 better than comparably sized CLIP models from OpenAI Radford et al. (2019) and LAION Schuhmann et al. (2021) respectively. Our DINOv2-ARL model demonstrates competitive performance across various datasets compared to LAION and OpenAI CLIP models. The relative performance of these models varies depending on the specific dataset. For example, on the Stanford Cars dataset, LAION-400m Schuhmann et al. (2021) CLIP outperforms OpenAI CLIP by a significant margin of over 12%. Conversely, for the Aircrafts dataset, both OpenAI CLIP and our DINOv2-ARL model show superior performance compared to LAION-400m CLIP. We believe this to be due to the differences in concept coverage for these particular datasets between the LAION400m, OpenAI WIT, and our MIX-CLASS-Collected datasets.
In zero-shot text retrieval, our model slightly outperforms or matches the next best CLIP model, LAION400M-CLIP VIT-L, with scores of 87.5 vs 87.6 on Flickr and 59.7 vs 60.1 on COCO. For image retrieval, our models show a significant advantage, achieving scores of 74.1 vs 70.2 on Flickr and 45.1 vs 43.0 on COCO. This improvement is likely due to the superior quality of the unimodal features produced by the DINOv2 and All-Roberta-Large-v1 encoders, compared to those of the multi-modal vision and text embeddings in the CLIP models.
5.2 0-shot Localization
One key advantage of leveraging frozen unimodal vision and text encoders is the enhancement provided by unimodal features. Specifically, the DINOv2 vision encoder’s robust localization capabilities enhance the joint embedding space of the DINOv2-ARL model when trained solely with projectors. We assess this through zero-shot segmentation performance, similar to the Bica et al. ; Mukhoti et al. (2023), as shown in Table 5. Our approach involves computing cosine similarities between each patch and all the ground truth classes and subsequently upscaling to the target size. Each patch is then classified into a corresponding class. Consistent with previous studies, the intersection over union (IoU) is computed solely for the foreground classes. In the zero-shot segmentation process of CLIP models, we employ a technique similar to Zhou et al. (2022) to alleviate the opposite visualization problem in CLIP models Li et al. (2023). The patch embeddings from the penultimate layer are passed through the value layer and output MLP of the final self-attention block, followed by projection into the joint embedding space using the vision projector. Meanwhile, our DINOv2-ARL model considers patch embeddings projected into the joint embedding space by the patch projector and augments them with the projected CLS token in a residual manner.
Our DINOv2-ARL model demonstrates superior performance compared to jointly trained dual encoder models like OpenAI’s CLIP, achieving over 8% improvement on Pascal VOC and over 10% on Pascal Context. Notably, models utilizing a fine-grained alignment loss like SPARC Bica et al. show improvements over CLIP. However, our DINOv2-ARL model outperforms SPARC by 4% on VOC and 3% on Context datasets. This underscores that the strong localization abilities of DINOv2 patch embeddings are retained even without training with a fine-grained alignment loss. We hypothesize that the localization performance could benefit from the quality of patch embeddings and a more precise localization alignment. Exploring fine-grained losses like SPARC with projector-only CLIP models presents an exciting direction for enhancing localization capabilities in VLMs.
@@ 5.3 Multi-Lingual Results
Our framework supports flexible swapping of text encoders, enabling multi-lingual capabilities through multi-lingual encoders, particularly beneficial for low-resource languages. We demonstrate this by aligning DINOv2-Large with paraphrase-multilingual-v2, chosen for its high CKA compatibility, using only English image-caption pairs. We then evaluated our model’s performance on multi-lingual image retrieval using the XTD dataset Aggarwal & Kale (2020b) and classification using the ImageNet dataset. For classification, we translated VDT prompts to the languages being considered using the nllb-200-distilled-600M Costa-jussà et al. (2022) model and applied them uniformly across all models.
Multi-lingual classification and retrieval results for five representative languages, are presented in Table 6. Detailed results are in Tables A.4, A.3. The lower section lists models trained exclusively with English captions, specifically the CLIP-VIT-L models trained on the WIT dataset Radford et al. (2021) and the LAION400M dataset Schuhmann et al. (2021). The upper sections feature models trained with translated captions, such as CLIP models based on LAION5B Schuhmann et al. (2022), M-CLIP models Chen et al. (2023a), and NLLB-CLIP models Visheratin (2023).
Our DINOv2-MpNet, trained solely on English image-caption pairs, outperforms other English-only CLIP models by over 31% in average retrieval performance across five languages and by 6% in English. While English CLIP models perform well on Latin script languages, their performance drops for non-Latin languages like RU and JP due to the English-only tokenizer. In contrast, our DINOv2-MpNet remains competitive across both Latin and non-Latin languages, even against models trained on multilingual data. Notably, it outperforms the laion5b-trained xlm-roberta-base-VitB32 by 0.6%, despite using only 20 million English image-caption pairs compared to the 5B multilingual pairs in LAION5B. In classification tasks, DINOv2-MpNet surpasses the LAION400m-trained ViT-L on English Imagenet, delivering significantly better results (over 20% on average) across five languages. Among multilingual models, it exceeds both nllb-clip and M-CLIP models, surpassing the next best M-CLIP/XLM-Roberta-Large-Vit-L-14 by over 8%, despite not using any multilingual text data. It also outperforms the LAION5B-trained CLIP model by 7% despite its use of multilingual image-caption pairs. This underscores the efficiency of our training approach, achieving highly performant models with significantly fewer image-caption pairs, and suggests that further training on translated pairs could enhance DINOv2-MpNet’s performance, particularly in low-resource languages.
@@ 5.4 Densely Captioned Images (DCI) Dataset and Long-Text Retrieval
The Densely Captioned Images (DCI) dataset Urbanek et al. (2024) offers a unique approach to image-text datasets, featuring 7,805 natural images with richly annotated, mask-aligned descriptions averaging over 1,000 words per image. This level of detail provides an opportunity to explore the limits of vision-language models in processing long-term textual information in relation to visual content. While DCI includes its own benchmarks using summarized captions, our focus is on image-text and text-image retrieval tasks using the entire dataset without summarization or subcropping, allowing us to investigate the long-text retrieval capabilities of our framework.
To demonstrate the advantages of processing longer captions, we conducted an experiment varying the maximum token length allowed by the tokenizer. As shown in Figure 7, our DINOv2-ARL encoder pair achieves comparable performance to OpenAI CLIP at the standard limit of 77 tokens. However, our approach’s strength becomes evident as we extend beyond this limit, with consistent improvement in retrieval accuracy up to approximately 200-300 tokens. These results highlight our framework’s ability to effectively utilize longer, more detailed captions for improved retrieval, capturing nuanced details and context that may be lost when constrained to shorter text sequences.
5.5 Training Compute
We report the Alignment Training compute requirements for different models in 7. We see that aligning pre-trained vision, language encoders to get a competitive CLIP like model requires only 50 hours of training with 8 A100 GPUS which is almost a 65 fold reduction in the amount of training compute. This makes the development of multi-modal models accessible to the wider research community as well as reducing the environmental impact of training highly performant multi-modal models by reusing strong publicly available uni-modal models. Since we only need to train 11.5M of the total 670M parameters (about 1 ) we can train with a much smaller and denser dataset reducing the data requirements to 20M which is 20 fold decrease in dataset requirement compared to CLIP models from LAION and OpenAI making our framework useful for training performant multi-modal models in various domains like mutli-modal systems for low-resource languages, 3D model search systems, fMRI to Image model mapping systems and many more. Despite the reduced compute and data requirements for alignment our model outperforms both CLIP models compared on domain transfer to Imagenet as well as image, text retrieval.
@@ 5.5 Training Compute
6 Related Works
Multimodal Pretraining: The CLIP models from OpenAI Radford et al. (2021) and ALIGN Jia et al. (2021) pioneered using web-scale image-caption data to align image and text modalities via an InfoNCE Oord et al. (2018) loss, optimizing mutual information between embeddings. LAION Schuhmann et al. (2021; 2022) replicated this approach in the open domain, open-sourcing pre-training datasets. While these models excel in zero-shot tasks, they demand substantial computational resources, around 20k GPU hours. Taking advantage of the recent improvements in the representation quality of unimodal encoders such as DINOv2 Oquab et al. (2023) (vision) and Sentence Transformer Reimers & Gurevych (2019) (language) models, Zhai et al. (2022) reduce the training cost by locking the image encoder and training only the text encoder to achieve competitive performance. Similarly, Khan & Fu (2023) further aligned frozen uni-modal encoders using projection layers, BitFit Zaken et al. (2021), and trainable adapters, but their approach is sub-optimal compared to CLIP, likely due to smaller datasets used and random encoder pair selection. In contrast, in this work, we strive to identify the best encoder pairs for alignment first and then scale up projector-only training to improve the multimodal alignment.
Representational Similarity: Recent studies show that the semantic similarity between vision and language model embeddings is high for several model pairs. Maniparambil et al. (2024) reports that this similarity, measured by Centered Kernel Alignment Kornblith et al. (2019), increases with more training data for vision models. Similarly, Huh et al. (2024) finds that better-performing language models have higher semantic similarity to the DINOv2 Oquab et al. (2023) vision model. These similarities have been leveraged for 0-shot and multi-lingual retrieval tasks using strong uni-modal encoders without additional training Maniparambil et al. (2024); Moschella et al. (2022), though scalability is an issue. Additionally, Merullo et al. (2022) demonstrates that a simple linear mapping allows a frozen language model to interpret visual input, provided the visual encoder aligns with language concepts (e.g., CLIP). Similarly, Dwivedi & Roig (2019); Dwivedi et al. (2020) also uses representational similarity metrics to identify pre-trained models for effective transfer to downstream tasks. These findings suggest that a simple projection transformation separates the embedding spaces of well-trained vision and language models, motivating our work on developing CLIP models using projection layers between semantically similar encoder pairs.
Automatic Data Curation: Our dataset curation pipeline draws on various approaches in Vision-Language dataset construction Radford et al. (2021); Gadre et al. (2024); Xu et al. (2024). Radford et al. (2021) used image metadata to gather high-quality image-caption pairs, while Schuhmann et al. (2021) replicated the CLIP dataset by filtering with pretrained vision encoders. Recent methods like Gadre et al. (2024) employ CLIP-based filtering and ad hoc filtering techniques, and Xu et al. (2024) mimics CLIP’s data collection via metadata retrieval. Similarly, Oquab et al. (2023) uses a pretrained vision encoder to curate web images most similar to images in curated datasets. Our approach is similar, constructing concept image prototypes from few-shot labeled examples and retrieving relevant web images from the LAION-400M pool using CLIP caption embeddings, avoiding the computational cost of generating vision embeddings for the entire dataset.
@@ 7 Conclusion
Our research introduces a paradigm shift in vision-language alignment, demonstrating that state-of-the-art performance can be achieved with a fraction of the resources traditionally required. By leveraging the latent compatibility of well-trained unimodal encoders, we have unlocked a new direction in efficient multimodal AI development.
Future work in this area could explore fine-grained alignment techniques, optimize projection architectures, and expand to other modalities beyond vision and language. By democratizing multimodal AI research, our framework has the potential to accelerate innovation and reshape approaches to multimodal AI development.
@@ Appendix A Appendix
A.1 CKA Preliminary
Centered Kernel Alignment (CKA) Kornblith et al. (2019) measures the similarity of induced graphs of concepts in each encoder space and can act as a guide for encoder pairs selection that are amenable to alignment as we demonstrate in section 4. We define CKA as follows: Given two sets of vectors and , CKA measures the similarity of these vectors in their respective high-dimensional feature spaces. The kernel matrices and are derived from the data sets and , respectively, and represent the inner products between the vectors in these spaces. The entries of and are computed as:
The CKA coefficient, , is defined as:
where HSIC stands for Hilbert-Schmidt Independence Criterion Gretton et al. (2005); Ma et al. (2020), which measures the dependence between the sets of vectors. This measure is invariant to orthogonal transformations and isotropic scaling of the data, making it robust for comparing different models.
A.2 CKA vs Graph structure

### v2

@@ 5.2 Impact of Projector Architectures
We ablate our projector combinations (1) for the DINOv2 and All-Roberta-Large-v1 encoders by training the projectors to convergence on the LAION-Class-Collected dataset and evaluating the performance on ImageNet 0-shot domain transfer. An MLP applied solely to the local vision tokens achieved 68.81% accuracy, while a Token projection [37] performed slightly better. Therefore, we used the Token projector for all tokens, both visual and textual. Adding projectors to the text side, targeting both text tokens and a global projector on the averaged local tokens (rows 3, 4, and 5), resulted in performance improvements. These projectors help transform the unimodal text encoder’s language-only representations to be more similar to the visual representations. Introducing projectors to the CLS token (row 6) of the visual encoder led to a significant performance increase from 72.15% to 75.13%. Using both CLS and patch projectors in tandem yielded the best performance at 76.12%. This improvement is attributed to DINOv2’s dual training objectives: the image-level DINO [7] objective on the CLS token and the patch-level iBOT [65] objective on the patch tokens learning effective global and local features.
@@ 5.3 Impact of Class-Collected Data / Retrieval Data
In Table 2, we ablate the different components of our alignment data. Specifically, we compare the high concept coverage LAION-CLASS-Collected dataset with the high semantic alignment retrieval datasets: CC3M, CC12M, and SBU. Our experiments show that aligning DINOv2 and All-Roberta-Large-v1 on the high concept coverage dataset results in a high ImageNet zero-shot domain transfer accuracy of 76.1 , though the retrieval accuracies are lower, at 52.7/42.2 due to the noisy semantic alignment in LAION dataset. In contrast, training with the higher image-caption quality retrieval datasets results in high image and text retrieval scores on the Flickr30k val set (85.3 and 72.4, respectively). However, the limited concept coverage of these datasets leads to a lower ImageNet accuracy of 54.1. Combining both types of datasets yields both high ImageNet accuracy and high image/text retrieval accuracies verifying that both dense coverage of unimodal spaces as well as high cross-modal semantic similarity is required to train effective projectors. To ensure that the extra data is adequately utilized, we train for an additional 15 epochs resulting in our best-performing model, achieving an ImageNet accuracy of 76.30 and Flickr retrieval scores of 87.54/74.17 (last row).
@@ 6 Results
We evaluate the alignment between vision and text encoders across commonly used VLM benchmarks, including zero-shot domain transfer, image retrieval 6.1, localization 6.2, multilingual classification/retrieval 6.3, and dense caption image-text retrieval 6.4. Our goal here is to evaluate the effectiveness of the learned alignment, showcase the flexibility of the framework as well as show that strong task-specific capabilities of uni-modal embeddings are retained in the joint embedding space. We demonstrate that aligning unimodal vision-language encoders can match or exceed the performance of large CLIP models, despite using smaller datasets and less compute. Additionally, our alignment framework is flexible, enabling the use of specialized encoders for specific tasks, such as aligning multilingual text encoders for multilingual or low-resource image classification/retrieval, or long-context text encoders for dense image/caption retrieval. Furthermore, aligning DINOv2 with a text encoder improves image localization beyond CLIP’s vision encoder due to DINOv2’s superior localization features.
@@
In this section we aim to evaluate the effectiveness of simple projection transformations in learning an alignment between semantically similar embedding spaces. Tables 3 and 4 report our model’s performance on zero-shot domain transfer to image classification datasets and image-text retrieval on the Flickr30k and COCO datasets, respectively. Detailed descriptions of the evaluation datasets can be found in the A.14, highlighting dataset domains, sizes, and prompt descriptions. We see that despite being trained on a 20M dataset our DINOv2-ARL projector model achieves an ImageNet accuracy of 76.3 which is 1 and 3.6 better than comparably sized CLIP models from OpenAI [45] and LAION [51] respectively. Our DINOv2-ARL model demonstrates competitive performance across various datasets compared to LAION and OpenAI CLIP models . The relative performance of these models varies depending on the specific dataset. For example, on the Stanford Cars dataset, LAION-400m [51] CLIP outperforms OpenAI CLIP by a significant margin of over 12%. Conversely, for the Aircrafts dataset, both OpenAI CLIP and our DINOv2-ARL model show superior performance compared to LAION-400m CLIP. We believe this to be due to the differences in concept coverage for these particular datasets between the LAION400m, OpenAI WIT, and our MIX-CLASS-Collected datasets.
In text retrieval, our model outperforms or matches the next best CLIP model, LAION400M-CLIP VIT-L, with scores of 87.5 vs 87.6 on Flickr and 59.7 vs 60.1 on COCO. For image retrieval, our models show a significant advantage, achieving scores of 74.1 vs 70.2 on Flickr and 45.1 vs 43.0 on COCO. This improvement is likely due to the superior quality of the unimodal features produced by the DINOv2 and All-Roberta-Large-v1 encoders, compared to those of the multi-modal vision and text embeddings in the CLIP models. These results demonstrate that simple projector transformations between uni-modal encoders can achieve competitive performance similar to models trained from scratch, providing further evidence that simple projection transformations separate semantically similar embedding spaces.
6.2 0-shot Localization
One key advantage of leveraging frozen unimodal vision and text encoders is the enhancement provided by unimodal features. Specifically, the DINOv2 vision encoder’s robust localization capabilities enhance the joint embedding space of the DINOv2-ARL model when trained solely with projectors. We assess this through zero-shot segmentation performance, similar to the [5, 37], as shown in Table 5. Our approach involves computing cosine similarities between each patch and all the ground truth classes and subsequently upscaling similarity maps to the target size. Each patch is then classified into a corresponding class. Consistent with previous studies, the intersection over union (IoU) is computed solely for the foreground classes.(Refer to Sec. A.10 for implementation details)
@@
Our DINOv2-ARL model demonstrates superior performance compared to jointly trained dual encoder models like OpenAI’s CLIP, achieving over 8% improvement on Pascal VOC and over 10% on Pascal Context. Notably, models utilizing a fine-grained alignment loss like SPARC [5] show improvements over CLIP. However, our DINOv2-ARL model outperforms SPARC by 4% on VOC and 3% on Context datasets. This underscores that the strong localization abilities of DINOv2 patch embeddings are retained even without training with a fine-grained alignment loss. We hypothesize that the localization performance could also benefit from a more precise localization alignment. Exploring fine-grained losses like SPARC with projector-only models presents an exciting direction for enhancing localization capabilities in VLMs. Additionally, the lower data demands of projector training may allow for the effective use of high-quality, smaller-scale grounding datasets to achieve precise alignment between word tokens and image patches in a supervised manner.
@@ 6.3 Multi-Lingual Results
Similar to the previous section, here we assess whether multi-lingual capabilities of a language encoder is retained when aligned to a vision encoder using projectors. We demonstrate this by aligning DINOv2-Large with paraphrase-multilingual-MpNetv2 (referred to as MpNet), chosen for its high CKA compatibility, using only English image-caption pairs and evaluating model performance on multi-lingual image retrieval on the XTD dataset [1] and classification on the ImageNet dataset. For classification, we translated the prompts to the considered languages using nllb-200-distilled-600M [12]. Multi-lingual classification and retrieval results for five representative languages are presented in Table 6 (For Detailed results Refer to Sec A.11 ) .The lower section lists models trained exclusively with English captions, [45] [51] while upper sections feature models trained with multi-lingual captions [50],[9], [58].
Our DINOv2-MpNet, trained solely on English image-caption pairs, outperforms other English-only CLIP models by over 31% in average retrieval performance across five languages and by 6% in English. DINOv2-MpNet remains competitive across both Latin and non-Latin languages, even against models trained on multilingual data. Notably, it outperforms the LAION5B trained xlm-roberta-base-VitB32 by 0.6%, despite using only 20 million English image-caption pairs compared to over 2B non-English pairs in LAION5B. A similar trend is observed in classification, with DINOv2-MpNet surpassing the next best English-trained model, by over 20% on average across five languages. Among multilingual models, the next best M-CLIP/XLM-Roberta-Large-Vit-L-14 by over 8%, despite not using any multilingual text data. DINOv2-MpNet’s robust multilingual performance, achieved without multilingual training data, demonstrates that MpNet’s capabilities are preserved in the joint embedding space through effective projector training of unimodal models.
@@ 6.4 Densely Captioned Images (DCI) Dataset and Long-Text Retrieval
We assess whether the ARL model maintains its long-context capabilities in the joint embedding space by conducting image and long caption retrieval on the Densely Captioned Images (DCI) dataset [57], which features caption pairs averaging over 1,000 words. Unlike DCI’s benchmarks that use summarized captions (see A.13), we focus on full image-text and text-image retrieval tasks without summarization or subcropping, enabling a comprehensive evaluation of our framework’s long-text retrieval capabilities.
To demonstrate the retention of long-context ability, we conducted an experiment varying the maximum token length allowed by the tokenizer. As shown in Figure 5, our DINOv2-ARL encoder pair achieves comparable performance to OpenAI CLIP at the standard limit of 77 tokens. However, our approach’s strength becomes evident as we extend beyond this limit, with consistent improvement in retrieval accuracy up to approximately 200-300 tokens. Given that DINOv2-ARL was trained with short-context image-caption pairs, these results underscore the model’s ability to retain long-context capabilities in the aligned joint embedding space.
6.5 Alignment Compute
We report the Alignment Training compute requirements for different models in 7. We see that aligning pre-trained vision, language encoders to get a competitive CLIP like model requires only 50 hours of training with 8 A100 GPUS which is almost a 65 fold reduction in the amount of alignment compute. This makes the development of multi-modal models accessible to the wider research community as well as reducing the environmental impact of training highly performant multi-modal models by reusing strong publicly available uni-modal models. Since we only need to train 11.5M of the total 670M parameters (about 1 ) we can train with a much smaller and denser dataset reducing the data requirements to 20M which is 20 fold decrease in dataset requirement compared to CLIP models from LAION and OpenAI making our framework useful for training performant multi-modal models in various domains like multi-modal systems for low-resource languages, 3D model search systems, fMRI to Image model mapping systems and many more where paired data is limited. Despite the reduced compute and data requirements for alignment, our model outperforms both CLIP models compared on domain transfer to Imagenet as well as image, text retrieval. Caveat: Our alignment assumes strong unimodal encoders are available and does not account for training compute. For completeness, DINOv2 was trained with 22k GPU (A100) hours, while ARL and MpNet used 7 TPUv3-8 for 400k steps [48].
@@ 7 Conclusion
Our research presents a significant advancement in vision-language alignment, showing that high performance can be attained with considerably fewer resources than usually needed. By utilizing the inherent compatibility of well-trained unimodal encoders, we offer a new perspective on efficient multimodal AI development.
Future efforts could investigate how our models might be integrated with Large Language Models, employ fine-grained alignment techniques, utilize different projection architectures, and extend to additional modalities beyond vision and language. Our framework may facilitate more accessible multimodal AI research, potentially speeding up innovation and influencing future approaches to multimodal AI development.
@@ Appendix A Appendix
A.1 Unlocking parts of text and vision encoders
We evaluated our model with different parts of the vision and text encoders unlocked for DinoV2-ARL, shown in Tab.A.3. Similar to Lit [63] we find that unlocking the vision encoder (e.g., via BitFit [61]) reduced performance, while full unlock resulted in unstable training. In contrast, unlocking the text encoder or applying slightly improved performance with increased training costs.
A.2 Training CLIP with same dataset
We compare our approach against CLIP-ViT-L models trained from scratch, and projector-only trained in Tab. A.3. We see that our 20M dataset is not enough to train the CLIP model (427M params) from scratch. Meanwhile, projector-only training of CLIP improves over OpenAI CLIP on COCO I2T and achieves competitive performance on Imagenet. Notably, none of the trained CLIP models outperform DINOv2-ARL.
A.3 Multi-lingual 0-shot Semantic Segmentation
The lower compute and paired data requirements of the framework lead to application flexibility simply by swapping the unimodal encoders. (see Sec. 6.2-6.4 in the main paper). An additional advantage of this flexibility is showcased in Fig. A.1 and Tab. A.3, where we use our aligned DINOv2-MpNet to perform multi-lingual semantic segmentation. Our segmentation scores stay consistent with different languages while CLIP often fails on non-english languages.
| Language | CLIP | DINOv2-MpNet |
| EN | 23.46 | 29.07 |
| ES | 18.86 | 28.69 |
| ZH | 8.46 | 28.06 |
| FR | 15.12 | 28.48 |
| DE | 21.30 | 27.91 |
| RU | 5.72 | 26.85 |
| Method (15 epochs) | Imagenet | COCO I2T |
| 67.67 | 53.16 | |
| 74.58 | 56.72 | |
| Text unlock | 75.90 | 56.62 |
| Projectors | 75.04 | 56.32 |
| Method (30 epochs) | Imagenet | COCO I2T |
| 50.30 | 36.12 | |
| 75.32 | 56.31 | |
| 72.10 | 59.04 | |
| DINOv2-ARL | 76.45 | 60.14 |
A.4 Toy Example using Random Latent Model
Similar to Sec. 3.2 here we investigate whether semantically similar encoder embedding spaces can be aligned through a simple projection transformation, using a random latent model.
In our experiment, we generated instances of two vector sets, and , each containing 32 vectors of 16 dimensions. Following the approach in [32, 21], we modeled the world using a latent distribution , with Image and Text representations ( and ) as random independent non-linear transformations from with additive noise. For each sampled pair of and matrices, we calculated the CKA and the minimum CLIP loss. The non-linear transform was defined as a randomly initialized 2-layer MLP with ReLU non-linearity and hidden dimensions significantly larger than the input dimensions, ensuring it could universally approximate the non-linear transformation [20]. Figure A.3 was used to generate each instance.
Figure A.3 illustrates the results of this experiment, showing a clear negative correlation between CKA and minima of the CLIP loss. As CKA increases, indicating greater similarity between the similarity structures of A and B, the minima of CLIP loss consistently decreases. Despite arising from a simplified experiment, the observed strong inverse relationship between CKA and CLIP loss provides empirical support for using CKA as a predictor of alignment potential between embedding spaces. Since CLIP loss is lower-bounded by mutual information, and mutual information is correlated with HSIC, higher CKA suggests a stronger alignment between embeddings. This implies that the achievable minima of CLIP loss is lower when the embedding spaces already have a higher CKA, reflecting greater mutual information and ease of alignment.
A.5 Embedding Graph structures visualized